LM3429: TOF CAMERA's LED DRIVER Schematic Problem Review

Part Number: LM3429

HI Subhash Chandra Sadhu

Test environment

OPT8241-CDK-EVM kit is used and IR LED circuit is in use Refer to attached circuit

OPT8241-CDK-EVM Using the firmware for the built-in laser and adjusting the Voxel SDK parameter without firmware modification.

- OPT8241-CDK-EVM OPT9221 ILLUM_N Signal line R99 1K Installed R101 does not use open processing ILLUM_P signal

LED DRIVER Schematic Problem Review

1. Circuits PL7 and ZD1 do not use LED DRIVER but 20V as power supply separately. The LED is driven without any problem.

    However, when 20V is input using LM3429 LED DRIVER, the switching noise is heard and the IR_LED_OUT voltage shakes.

Thank you Subhash Chandra Sadhu6330.tof_sch_170327.pdf

 

17 Replies

  • I am trying to get one of these two identical threads moved to the "optical sensors" forum as nobody in this forum is familiar with that EVM. Thanks.
  • In reply to Clinton Jensen:

    TOF CAMERA's LED DRIVER problem should be solved. Please let me know your engineer name

    Clinton Jensen Thank you

  • Hi!

    I have taken a cursory look at your schematic. I believe that you have a problem of the output voltage of the boost converter LM3429 dropping out of regulation whenever the camera is being driven.
    Can you confirm that you can actually hear the PCB make some sound? Can you try increasing the register value of pix_cnt_max to 1.5 times the original value and see if the sound reduces in its frequency?

    Regards,
    Subhash

    3D ToF Product Landing Page

  • In reply to Subhash Chandra Sadhu:

    hi Subhash Chandra Sadhu

    - You said the boost converter output voltage is a problem. Which part is the problem?

      Led directly to IR_LED_VOUT. Then the output voltage is not a problem.

    - Please look at the sending circuit again. And answer

     And if you connect the ILLUM_EN signal from the tof camera to Figure 2, you hear the switching noise and the output voltage shakes.

     tof_sch_170329.pdfWhat is the problem

  • In reply to MINSU PARK:

    Hi!

    Given your schematic I am guessing that you have an issue of LM3429 falling out of regulation. A potential reason could be that the input supply to LM3429 can't sustain the power requirement. Can you compare the IC power requirement to the power source capability?
    Also, we expect the output voltage of LM3429 to reduce when LM3429 is disabled by the ILLUM_ENz signal. So that is an expected behavior of the IC.
    When you say you hear switching noise, can you confirm if you actually hear sound from the PCB? Highly varying voltages on the power supply can cause the capacitors to generate some sound because of mild piezoelectric nature of the dielectric in the capacitors.

    Regards,
    Subhash

    3D ToF Product Landing Page

  • In reply to Subhash Chandra Sadhu:

    hi Subhash
    Thanks for the answer.
    - The pix_cnt_max parameter value was changed and the switching noise was reduced. What is the cause
    - The input power of the lm3429 is 5V
    - Switching sound seems to be heard on PL8 coil parts
    Thanks ,Subhash
  • In reply to MINSU PARK:

    Hi!

    The cause of the noise is a combination piezoelectric nature of the ceramic capacitors and magnetostrictive effect of the inductor. This is a well known problem. Please refer to the following links:
    - www.capacitorindustry.com/.../
    - en.wikipedia.org/.../Coil_noise

    Can you verify if your 5V supply can give enough current for the peak power requirement of LM3429?

    Do you see if the camera image is looking bad? If not, your only problem may be that your board makes some acoustic noise. If you see excessive flickering in your camera image, you may have to do further debug.

    Regards,
    Subhash

    3D ToF Product Landing Page

  • In reply to Subhash Chandra Sadhu:

    hi Subhash

    Thanks for the answer.

    - 3A power supply to give the LM3429 5V power. The LM3429 has enough current.

     Also, as you can see from the circuit diagram, there is no problem in driving the LED.

    - The camera image does not look bad. Switching noise sound is a problem.

    - I connected the PWM signal applied to the LM3429 NDIM pin to the beacon instead of the ILLUM_EN signal. There is no switching noise and the LED works well.

    What is the cause

    Thanks

  • In reply to MINSU PARK:

    hi Subhash

    - The OPT8241-CDK-EVM firmware is using firmware for the laser. Is there firmware that uses the ir LED applied to the OPT8241-CDK-EVM kit?

    - If you have the firmware for the LEDs applied to the OPT8241-CDK-EVM kit, can you send it to me

    Thanks

  • In reply to MINSU PARK:

    Hi!

    The OPT8241 chipset is reconfigurable to work in multiple modes of operation. To make the LEDs work, you would have to keep the modulation frequency at a value less than 20MHz. The LEDs could dissipate a lot of heat. So I would recommend that you start using the board at a relatively low (4%) integration duty cycles. Once you get the image and you are sure that the board is not overheating, you could carefully increase the integration duty cycle to improve depth performance.
    These changes are possible by just modifying register settings in Voxel viewer. A firmware change is not required for these changes.

    Regards,
    Subhash

    3D ToF Product Landing Page